# The BRAVO Semantic Segmentation Challenge Results in UNCV2024

### v1

@@ 2.2 Track 1: DINOv2-OOD – Eindhoven University of Technology
DINOv2, ViT-B, 16x16 patch size, linear decoder 75.5 70.5 81.6
@@
DeiT III (IN1K), ViT-S, 16x16 patch size, linear decoder 59.4 62.8 47.6
@@
Results.  As shown in Tab. 3. Our best performing model, DINOv2 with a ViT-L/8 backbone and a linear decoder, achieves the highest BRAVO index of $77.9$, which is $+10.1$ higher than the next best method, i.e., PixelOOD presented in Sec. 2.3.
@@
DINOv2, ViT-L, 8x8 patch size, linear decoder 76.7 40.0 99.4 91.4 2.1 38.2
DINOv2, ViT-L, 16x16 patch size, linear decoder 75.9 41.2 99.5 92.3 1.8 37.2
DINOv2, ViT-g, 16x16 patch size, linear decoder 75.9 39.3 99.5 91.8 1.8 37.1
DINOv2, ViT-B, 16x16 patch size, linear decoder 74.5 43.3 99.4 92.3 2.1 39.5
DINOv2, ViT-S, 16x16 patch size, linear decoder 66.9 45.1 99.1 92.5 2.5 43.8
DINOv2, ViT-g, 16x16 patch size, Mask2Former decoder 49.1 23.2 98.6 81.4 5.1 79.4
PixOOD [26] 73.6 44.8 99.1 91.8 10.0 38.3
Ensemble C 43.2 58.5 93.5 84.0 15.5 55.4
DeiT III (IN1K), ViT-S, 16x16 patch size, linear decoder 64.8 54.7 97.9 89.2 1.8 52.8
Baseline: SegFormer-B5 [30] 57.4 24.1 98.3 77.2 4.1 71.5
Baseline: ObsNet-R101-DLv3plus [2] 67.3 34.1 98.2 87.1 4.0 51.5
Baseline: Mask2Former-SwinB [5] 28.6 13.2 97.4 47.0 55.2 82.3
@@
Another noteworthy observation is that all our models have relatively low ECE values, indicating that they are well-calibrated, even though no explicit calibration techniques were applied. Even the DeiT-III-based model, which scores low on the overall BRAVO score, achieves a low ECE value of 1.8. Compared to other models in the benchmark, all our models use the single-scale ViT architecture as the backbone. This might contribute to the good calibration of our models, but further investigation is needed to confirm this.
@@
OOD results are detailed in Tab. 5. Surprisingly, our configuration with worst confidence estimation for valid pixels, DINOv2 with ViT-g/16 and Mask2Former, achieves the highest AUPRC of 84.1, the highest AUROC of 98.8, and the lowest FPR@95 of 4.5 for detecting invalid pixels. This suggests that the mask classification framework used by Mask2Former, where per-class masks are predicted separately, allows this decoder to more accurately identify which pixels belong to the mask of an in-distribution class and which do not. Additionally, while a smaller patch size results in worse confidence estimation for valid pixels, it helps in identifying invalid pixels. Qualitative analyses show that the smaller patch size enables the model to better separate valid and invalid pixels, as it can capture more fine-grained details.
@@
DINOv2, ViT-B, 16x16 patch size, linear decoder 70.6 96.0 16.7
@@ 0.B.1 Track 1: DINOv2-OOD – Eindhoven University of Technology
Method.Fig. 3 overviews the DINOv2-OOD approach.
@@
DINOv2, ViT-g, 16x16 patch size, linear decoder 68.9 90.4 71.1 73.2 74.7 73.2
DINOv2, ViT-B, 16x16 patch size, linear decoder 66.5 87.9 71.2 72.8 74.0 73.0
DINOv2, ViT-S, 16x16 patch size, linear decoder 66.9 88.9 72.0 72.6 74.3 73.5
DINOv2, ViT-g, 16x16 patch size, Mask2Former decoder 49.1 66.7 49.0 53.9 49.9 50.0
DeiT III (IN1K), ViT-S, 16x16 patch size, linear decoder 47.0 62.1 48.4 65.1 47.6 56.4
@@ 0.B.2 Track 1: PixOOD – Czech Technical University in Prague
Method.Fig. 4 visualizes the three PixOOD variants that were submitted to the BRAVO challenge 2024.
@@ 0.B.3 Track 1: PhyFea – University Of Bologna & ETH Zurich
Method.Fig. 5 overviews the PhyFea approach.
@@ 0.B.4 Track 2: InternImage-OOD – CASIA & Objecteye
Method.Fig. 6 overviews the InternImage-OOD solution.

### v2

@@ 2.2 Track 1: DINOv2-OOD – Eindhoven University of Technology
DINOv2, ViT-B, 16x16 patch size, linear decoder 75.5 70.5 81.4
@@
DeiT III (IN1K), ViT-S, 16x16 patch size, linear decoder 54.1 62.8 47.6
@@
Results.  As shown in Tab. 3. Our best performing model, DINOv2 with a ViT-L/8 backbone and a linear decoder, achieves the highest BRAVO index of $77.9$, which is $+16.7$ higher than the next best method, i.e., PixelOOD presented in Sec. 2.3.
@@
DINOv2, ViT-L, 8x8 patch size, linear decoder 76.7 40.0 99.4 91.4 2.0 38.8
DINOv2, ViT-L, 16x16 patch size, linear decoder 75.9 41.2 99.5 92.3 1.7 37.8
DINOv2, ViT-g, 16x16 patch size, linear decoder 77.6 39.3 99.5 92.3 1.8 37.6
DINOv2, ViT-B, 16x16 patch size, linear decoder 71.7 43.3 99.4 92.3 2.1 40.3
DINOv2, ViT-S, 16x16 patch size, linear decoder 66.5 45.1 99.2 91.8 2.5 44.3
DINOv2, ViT-g, 16x16 patch size, Mask2Former decoder 78.2 23.2 99.2 87.9 5.0 63.6
PixOOD w/ ResNet-101 DeepLabv3 [26] 43.2 58.5 93.5 84.0 15.1 54.6
Ensemble C 73.9 47.4 99.1 92.5 52.7 34.7
DeiT III (IN1K), ViT-S, 16x16 patch size, linear decoder 44.9 54.7 97.9 89.2 1.7 54.2
Baseline: SegFormer-B5 [30] 67.4 24.1 97.2 77.2 30.7 71.9
Baseline: ObsNet-R101-DLv3plus [2] 65.3 32.1 98.5 87.8 45.6 63.4
Baseline: Mask2Former-SwinB [5] 67.2 13.2 90.4 47.0 55.1 82.8
@@
Another noteworthy observation is that all our models have relatively low ECE values, indicating that they are well-calibrated, even though no explicit calibration techniques were applied. Even the DeiT-III-based model, which scores low on the overall BRAVO score, achieves a low ECE value of 1.7. Therefore, further investigation is needed to understand why the ECE values are so low.
@@
OOD results are detailed in Tab. 5. Surprisingly, our configuration with worst confidence estimation for valid pixels, DINOv2 with ViT-g/16 and Mask2Former, achieves the highest AUPRC of 84.1, the highest AUROC of 98.8, and the lowest FPR@95 of 4.5 for detecting invalid pixels. This suggests that the mask classification framework used by Mask2Former, where per-class masks are predicted separately, allows this decoder to more accurately identify which pixels belong to the mask of an in-distribution class and which do not. Additionally, while a smaller patch size results in worse confidence estimation for valid pixels (see Fig. 3), it helps in identifying invalid pixels. Qualitative analyses show that the smaller patch size enables the model to better separate valid and invalid pixels, as it can capture more fine-grained details. Finally, while scaling from ViT-L to ViT-g improves mIoU for valid pixels (see Tab. 4), OOD detection performance shows a noticeable degradation.
@@
DINOv2, ViT-B, 16x16 patch size, linear decoder 70.6 96.6 15.1
@@ 2.3 Track 1: PixOOD – Czech Technical University in Prague
[…]
PixOOD 53.5 40.4 79.1
PixOOD w/ DeepLabv3 Decoder 59.4 46.1 83.5
PixOOD w/ ResNet-101 DeepLabv3 61.2 58.7 64.0
Table 6: Track 1 – PixOOD – Analysis.
@@ 0.B.1 Track 1: DINOv2-OOD – Eindhoven University of Technology
Method.Figure 3 overviews the DINOv2-OOD approach.
@@
DINOv2, ViT-g, 16x16 patch size, linear decoder 67.7 88.2 71.0 73.2 74.7 73.2
DINOv2, ViT-B, 16x16 patch size, linear decoder 68.5 87.9 71.2 72.8 74.0 73.0
DINOv2, ViT-S, 16x16 patch size, linear decoder 66.9 83.1 70.2 70.6 68.6 72.9
DINOv2, ViT-g, 16x16 patch size, Mask2Former decoder 49.1 94.4 40.9 53.9 64.3 60.1
DeiT III (IN1K), ViT-S, 16x16 patch size, linear decoder 58.8 50.1 65.1 71.4 58.5 65.6
@@ 0.B.2 Track 1: PixOOD – Czech Technical University in Prague
Method.Figure 4 visualizes the three PixOOD variants that were submitted to the BRAVO challenge 2024.
@@ 0.B.3 Track 1: PhyFea – University Of Bologna & ETH Zurich
Method.Figure 5 overviews the PhyFea approach.
@@ 0.B.4 Track 2: InternImage-OOD – CASIA & Objecteye
Method.Figure 6 overviews the InternImage-OOD solution.